# Deformable Mamba for Wide Field of View Segmentation

### v1

@@ Abstract
Wide-FoV cameras, like fisheye and panoramic setups, are essential for broader perception but introduce significant distortions in 180° and 360° images, complicating dense prediction tasks. For instance, existing MAMBA models lacking distortion-aware capacity cannot perform well in panoramic semantic segmentation. To address this problem, this work presents Deformable Mamba, a unified framework specifically designed to address imaging distortions within the context of panoramic and fisheye semantic segmentation. At the core is a decoder constructed with a series of Deformable Mamba Fusion (DMF) blocks, making the whole framework more deformable, efficient, and accurate, when handling extreme distortions. Extensive evaluations across five datasets demonstrate that our method consistently improves segmentation accuracy compared to the previous state-of-the-art methods tailored for specific FoVs. Notably, Deformable Mamba achieves a +2.5% performance improvement on the 360° Stanford2D3D dataset, and shows better results across FoVs from 60° to 360°.
Comparison of wide Field of View (wide FoV) segmentation.
Omni-directional performance across wide FoV from 60° to 360°.
Figure 1: Wide Field of View (wide-FoV) segmentation is investigated on 180° and 360° images. (a) Compared to other methods tailored for 90° or 180° data, our unified Deformable Mamba framework generalizes well on wide-FoV segmentation. (b) Besides, our methods yield consistently better results (mIoU %) across all different FoVs as compared to the baseline on the 360° Matterport3D dataset.
@@ 1 Introduction
Dense image analysis like image segmentation [2, 58, 21] is a fundamental task in computer vision, forming a crucial component of numerous downstream vision-based applications [70]. At the same time, advancements in sensor technology have led to a variety of sensor types, each with unique characteristics that impact imaging. In recent years, diverse sensors such as pinhole cameras, fisheye cameras [64, 48] with 180° field of views (FoV), and even panoramic cameras [3, 1, 67] with 360° FoV have been increasingly adopted across vision tasks. However, differences in camera construction and sensor type result in distinct imaging characteristics, which can complicate traditional image processing [66]. Adapting algorithms to perform effectively with a specific FoV often requires considerable time and resources, involving method redesign, parameter tuning, and model retraining to address the unique challenges posed by each camera type.
@@
Based on our experiments (Fig. Deformable Mamba for Wide Field of View Segmentation), we observe that straightforwardly applying or transferring existing specific models to new sensor types frequently leads to significant performance degradation. For instance, methods [21] specifically designed for narrow-FoV pinhole cameras typically cannot generalize well when used for analyzing fisheye images with 180° FoV. This is primarily due to the high degree of image distortion and object deformation present in wide-FoV cameras, which complicates model generalization. Standard methods struggle to accurately capture the diverse object shapes and spatial distributions caused by such deformations. For instance, as shown in Fig. 2, the altered pixel arrangement in wide-FoV images might lead to misclassifications and limits their adaptability. Previous methods have proposed solutions targeting specific sensor types, such as CNN-based and transformer-based semantic segmentation models [52, 31] optimized for fisheye images, as well as approaches [66, 67] tailored to mitigate distortion in panoramic segmentation. So far, to address diverse distortions across different camera types in a unified manner remains under-explored.
@@
Figure 2: Comparison of camera imaging with pinhole (left) and fisheye (right) cameras with different wide-FoV. Pinhole cameras maintain geometric fidelity but limited coverage, whereas wide-FoV cameras offer expansive scene capture while introducing substantial geometric distortions.
To address these challenges, we propose the Deformable Mamba model (DMamba for short), a versatile solution for image segmentation across varying wide-FoVs. In this work, a unified framework is created to effectively handling images from multiple camera types and sources, spanning FoVs from 180° to 360°, from indoor to outdoor scenarios, in synthetic and real-world imageries. At the core of our approach is a novel deformable decoder, which synergistically integrates State Space Models (SSMs) with deformable convolutions. Specifically, we leverage quadri-directional selective scan operations derived from VMamba [37] to capture global contextual information, while simultaneously employing deformable convolutions [77] to enhance feature representation. This dual-pathway desin effectively addresses the limitations of scanning mechanisms of Mamba [16] in visual applications while also endowing the model with the ability to perceive distortion. Our deformable decoder accommodates varied receptive fields, enabling seamless integration with CNN-based, transformer-based, and Mamba-based architectures. By leveraging this unified decoder, Deformable Mamba effectively captures features across disparate imaging perspectives, making it an adaptable solution for segmentation tasks across a broad spectrum of camera types.
To evaluate the effectiveness of Deformable Mamba, we conducted extensive experiments across five datasets for semantic segmentation, including three 360° (SynPASS [67], Stanford2D3D [1], Matterport3D [3]) and two 180° (WoodScape [64], SynWoodScape [48]) datasets. The data encompass a variety of scenarios including 180° and 360° images, indoor and outdoor environments, and both synthetic and real-world data. Our unified framework performs consistently well across various datasets. Notably, on 360° Stanford2D3D dataset, Deformable Mamba achieved state-of-the-art performance, with a ${+}2.5\%$ absolute improvement over the pinhole Mamba model. In 180° SynWoodScape dataset, it similarly surpassed prior best-performing models, achieving an increase of ${+}2.2\%$ in mean Intersection over Union (mIoU). Further ablation studies, including comprehensive omni-directional comparisons, consistently demonstrated the model’s effectiveness in handling wide-FoVs, showing its potential as a general solution for dense image analysis across diverse sensor types. We hope this work opens up new possibilities for unifying wide-FoV segmentation.
To summarize, our contributions are as follows.
  • We propose a Mamba-based framework termed Deformable Mamba, with the aim at unifying distortion-aware wide-FoV segmentation.
  • A deformable decoder constructed with Deformable Mamba Fusion (DMF) module is specifically designed to boost the performance of mamba-based models in the wide-FoV segmentation task.
@@ 2.1 Wide Field of View Segmentation
Wide Field-of-View (Wide-FoV) semantic segmentation has received increasing attention as it enables broader scene understanding in applications such as autonomous driving and robotics. However, directly applying pinhole models [39, 2, 5, 15, 26, 65, 65, 36, 34, 72, 50, 58, 9, 19, 6, 70] to wide-FoV images often results in downgraded performance due to object deformations and image distortions. Previous work [61, 62] propose the Panoramic Annular Semantic Segmentation (PASS) framework using a panoramic annular lens system.  [27] introduces panoramic panoptic segmentation for panoramic views. In [66, 67], deformable modules are proposed to enhance the distortion perception capability in transformers. Zheng et al. [71] introduce the Open Panoramic Segmentation (OPS) task in an open-vocabulary context. Apart from panoramas, previous methods [31, 11] revolve around multi-task on fisheye images, while a pretraining paradigm [42] and semi-supervised approaches [45, 43] are introduced to address the challenges of fisheye data. Manzoor et al. [40] substitutes the conventional convolutional modules in U-Net [47] and residual U-Net [46] with restricted deformable convolution (RDC) [11], imparting fisheye-distortion awareness to the model. Departing from these approaches, we are the first to propose a unified model applicable to all wide-FoV inputs, including panoramic and fisheye images.
2.2 State Space Model
CNNs [30] have achieved remarkable success in computer vision through effective local feature extraction and hierarchical representation learning capabilities, while Transformers [54] excel at modeling long-range dependencies despite their quadratic computational complexity. Recently, the state-space-model (SSM) based Mamba [16] architecture has emerged as a promising alternative, offering linear complexity and parallel computation through its selective scan mechanism, ushering in a new paradigm for long sequence modeling. The emergence of Mamba in sequence modeling has naturally extended to vision tasks, Zhu et al. [75] adopt the patch embedding from Vision Transformer (ViT) [12] to project 2D image patches into sequential tokens compatible with Mamba architecture. To address the inherent discrepancy between image and sequence representations, a quadri-directional scanning and merging mechanism termed SS2D is introduced in [37] to supersede Mamba’s canonical 1D scan operation. Subsequent work [25, 44, 24, 32, 69, 63, 4] introduce various scanning methods to explore variations of Mamba in visual representations. However, the study of using SSM for wide-FoV imagery is under-explored. Therefore, for the first time, we propose deformable Mamba to address both panoramic and fisheye semantic segmentation.
@@ 3 Methodology
3.1 Preliminary
Image Distortion. As shown in the left side of Fig. 2, the point $A$ in the world coordinates undergoes extrinsic transformation to camera coordinates by Eq. (1), where $R$ and $T$ denote the rotation matrix and translation vector, respectively. Following Eq. (2), points in the camera coordinates are projected onto the image plane, with $f_{x}$ and $f_{y}$ representing the focal lengths and $c_{x},c_{y}$ defining the principal point. While straight lines are preserved in pinhole-approximate narrow-FoV cameras, only points within the sensor’s FoV (red box in Fig. 2) are imaged, excluding points like $B^{\prime}$ that fall outside this region. Narrow-FoV cameras follow perspective projection principles where object size in the image varies inversely with distance between object and camera, introducing moderate distortion.
$\begin{bmatrix}\mathbf{X}_{c}\\ \mathbf{Y}_{c}\\ \mathbf{Z}_{c}\end{bmatrix}=\mathbf{R}\begin{bmatrix}\mathbf{X}_{w}\\ \mathbf{Y}_{w}\\ \mathbf{Z}_{w}\end{bmatrix}+\mathbf{T},$ (1)
$\begin{bmatrix}\mathbf{x^{\prime}}\\ \mathbf{y^{\prime}}\\ 1\end{bmatrix}=\begin{bmatrix}\mathbf{f}_{x}&0&\mathbf{c}_{x}\\ 0&\mathbf{f}_{y}&\mathbf{c}_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}\frac{\mathbf{X}_{c}}{\mathbf{Z}_{c}}\\ \frac{\mathbf{Y}_{c}}{\mathbf{Z}_{c}}\\ 1\end{bmatrix}.$ (2)
In contrast, wide-FoV cameras like fisheye cameras employ complex projection models involving light refraction. Kannala et al. [28] propose a generalized projection model for wide-FoV cameras, the distance between the image point and the principal point $r$ is approximated by:
$\mathbf{r}(\mathbf{\theta})=\mathbf{k_{1}}\mathbf{\theta}+\mathbf{k_{2}}\mathbf{\theta}^{3}+\mathbf{k_{3}}\mathbf{\theta}^{5}+\mathbf{k_{4}}\mathbf{\theta}^{7}+\mathbf{k_{5}}\mathbf{\theta}^{9},$ (3)
where $\theta$ is the angle between the principal axis and the incoming ray, $k_{i},i\in\{1,2,3,4,5\}$ are the parameters of wide-FoV camera model. While these cameras capture wider FoV, they suffer from more severe geometric distortions compared to the perspective projection of narrow-FoV cameras. As shown in the right side of Fig. 2, barrel distortion in wide-FoV cameras causes straight lines to appear curved outward, particularly pronounced near image edges while the center remains relatively undistorted. This distortion manifests as radially-increasing displacement from the center, resulting in progressive spatial compression toward the image periphery.
State Space Model. State space model (SSM), which describes the dynamics of a system, can be formulated as:
where $x(t)$ and $y(t)$ denote the inputs and outputs of the system, respectively, $h(t)$ is the inner states of the system, $A,B,C$ are the system parameters. Gu et al.  [18] propose Structured State Space Sequence Model (S4) by initializing the parameters $A$ using Hippo Matrix (5) and discretizing parameters A and B by step size $\Delta$.
$\mathbf{A}_{nk}=-\begin{cases}(2\mathbf{n}+1)^{1/2}(2\mathbf{k}+1)^{1/2}&\text{if }\mathbf{n}>\mathbf{k}\\ \mathbf{n}+1,&\text{if }\mathbf{n}=\mathbf{k}\\ 0,&\text{if }\mathbf{n}<\mathbf{k}.\end{cases}$ (5)
In Mamba [16], a selective scan mechanism is proposed to further extend the SSM from time-invariant to time-variant, while also enabling parallel training. Observations. Based on HiPPO theory [17], Mamba [16] is inherently a sequence model, where the output of the current token only depends on preceding tokens, without considering subsequent ones - a limitation that becomes problematic for 2D images with inherent spatial structures. Unlike Transformer’s [54] mechanism in ViT [12], where attention computation for each patch considers all patches globally, Mamba processes image patches sequentially.
To address this limitation, previous work [37, 75, 4, 76, 63] have explored different scanning patterns, including bidirectional or quadri-directional scanning approaches, partially addressing the challenges posed by Mamba’s unidirectional scanning nature in visual tasks. However, these multi-directional scanning approaches face another challenge. As illustrated in Fig 3., consider an input image embedded into 9 patches (labeled 1-9). When modeling patch 5 using SS2D’s [37] quadri-directional scanning mechanism, the scanning sequences are:
  • (1) Left-Right: 1 - 2 - 3 - 4 - 5
  • (2) Right-Left: 9 - 8 - 7 - 6 - 5
  • (3) Top-Down: 1 - 4 - 7 - 2 -5
  • (4) Down-Top: 9 - 6 - 3 - 8 - 5
Figure 3: Visualization of SS2D quadri-directional scanning.
Patch 1 and patch 3 maintain equal distances from patch 5, demonstrating uniform spatial distribution. As illustrated in sequences (1) (3) (4), the scanning trajectory from patch 1 to patch 5 traverses double the path length compared to the direct route between patch 3 and patch 5, despite their equivalent spatial distances. Since the Mamba model learns relationships between serialized tokens, the contribution of earlier patches in the scanning path to the current patch decreases as their distance from the current patch increases. Thus, along this scanning path, we infer that patch 1 is more important to patch 5 than patch 3. This priority is introduced by the scanning mechanism, but it is not a reasonable assumption. Additionally, when computing for other target patches, this priority changes based on the target’s position.
@@ 3.2 Overall Architecture
More specifically, given the wide-FoV input image in the shape of $H\times W\times 3$, a stem module is firstly utilized to patchify the original image into the shape of $\frac{H}{4}\times\frac{W}{4}\times 3$, the encoder gradually down-samples feature maps in the $l^{th}$ stage with channel dimensions $C_{l}\in\{96,192,384,768\}$ and resolutions $\in\{\frac{H}{4}\times\frac{W}{4},\frac{H}{8}\times\frac{W}{8},\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{H}{32}\}$. The features extracted by the encoder undergo a hierarchical fusion process in the decoder.
@@ 3.3 Deformable Mamba Decoder
The Deformable Mamba Decoder comprises four fusion and upsampling modules, where we introduce a novel Deformable Mamba Fusion (DMF) module for multi-scale feature integration. The initial DMF block processes dual inputs of ${\frac{H}{32}\times\frac{H}{32}\times 768}$ feature maps derived from the final encoder layer. The fused features are then processed through an upsampling module that simultaneously reduces channel dimensionality and increases spatial resolution, followed by channel mixing via an MLP layer. Following [49, 68, 35, 60], we utilize Pixel Shuffle [49] as the upsampling operation. In the subsequent stages, the second through fourth DMF blocks implement skip connections, facilitating interaction between the previously fused and upsampled features and their corresponding encoder features at matching scales. This progressive fusion strategy enables comprehensive multi-scale feature integration. The decoder’s output undergoes final processing through a fully connected layer and class mapping to generate the segmentation predictions.
3.4 Deformable Mamba Fusion
To address the aforementioned limitations of Mamba’s scanning mechanism in visual tasks (Sec. 3.1), we propose the DMF module, which simultaneously adapts the features extracted by the Mamba-based encoder and integrates multi-scale features. As illustrated in the lower part of Fig. 4, the DMF module takes two inputs: one is the feature $\mathbf{D}_{j}$ with dimensions $c\times h\times w$ output from the upper decoder layer, and the other is the corresponding encoder features $\mathbf{E}_{i}$ at matching scales, where $i$ and $j$ satisfy the relationship $i+j=5$. We employ the Cross-Scan [37] that scans $\mathbf{D}_{j}$ in four directions: (1) Left to Right, (2) Right to Left, (3) Top to Down, and (4) Down to Top, utilizing Mamba model’s global perception to generate four feature maps with resolution $c\times h\times w$ Following Cross-Merge [37], we first reverse features (2) and (4) along their last dimension and add them to (1) and (3) respectively, then transpose the $h$ and $w$ dimensions of feature (3) before adding it to (1).
Inspired by previous works [10, 77, 56, 59], we propose a simple yet effective approach, utilizing deformable convolution [77] to enhance the features extracted by the Mamba model. Specifically, for a feature point in the encoder feature map $\mathbf{E}_{i}$, we attend to $\mathbf{K}$ neighboring feature points, allowing the model to learn offsets that compensate for the biases introduced by the scanning process. Additionally, convolution operations on these selected regions provide a more localized attention mechanism, further enhancing the visual representation capability of Mamba-based models in visual tasks. The mathematical formulation is as follows:
$\mathbf{y}(\mathbf{p})=\sum^{K}_{k=1}\mathbf{w}_{k}\mathbf{m}_{k}\mathbf{x}(\mathbf{p}+\mathbf{p}_{k}+\Delta\mathbf{p}_{k}),$ (6)
@@ 3.4 Deformable Mamba Fusion
where $\mathbf{p}$ denotes the current location in feature map, $\mathbf{w_{k}}$ and $\mathbf{p_{k}}$ denote the weight and initial offsets for the $\mathbf{k}$-th location, respectively. And $\mathbf{m_{k}}$ is a learnable modulation scalar, which helps to modulate the input feature amplitudes from different spatial locations. For example, $\mathbf{K}=9$ and $\mathbf{p_{k}}\in{(-1,-1),...,(1,1)}$ defines a $\mathbf{3\times 3}$ convolutional kernel with dilation 1.
The outputs from these two branches are concatenated and fed into a convolutional layer for feature fusion. This simple yet effective architecture successfully adapts the features extracted by the Mamba model, efficiently integrates multi-scale features, while endowing the entire decoder with deformable capabilities. Experiments on multiple datasets demonstrate both the effectiveness and efficiency of our proposed model.

### v2

@@ Abstract
Recent advancements in the Mamba architecture, with its linear computational complexity, being a promising alternative to transformer architectures suffering from quadratic complexity. While existing works primarily focus on adapting Mamba as vision encoders, the critical role of task-specific Mamba decoders remains under-explored, particularly for distortion-prone dense prediction tasks. This paper addresses two interconnected challenges: (1) The design of a Mamba-based decoder that seamlessly adapts to various architectures (e.g., CNN-, Transformer-, and Mamba-based backbones), and (2) The performance degradation in decoders lacking distortion-aware capability when processing wide-FoV images (e.g., 180° fisheye and 360° panoramic settings). We propose the Deformable Mamba Decoder, an efficient distortion-aware decoder that integrates Mamba’s computational efficiency with adaptive distortion awareness. Comprehensive experiments on five wide-FoV segmentation benchmarks validate its effectiveness. Notably, our decoder achieves a +2.5% performance improvement on the 360° Stanford2D3D segmentation benchmark while reducing 72% parameters and 97% FLOPs, as compared to the widely-used decoder heads.
@@
(a) Comparison across narrow- and wide-FoV semantic segmentation.
(b) Performance and efficiency comparison of different decoders on Stanford2D3D [1] panoramic dataset with the same backbone [34].
Figure 1: Our deformable Mamba (a) achieves better results across wide-FoV semantic segmentation while (b) maintaining parameter and computational efficiency.
@@ 1 Introduction
Dense image analysis like image segmentation [2, 55, 20] is a fundamental task in computer vision, forming a crucial component of numerous downstream vision-based applications [67]. At the same time, advancements in sensor technology have led to a variety of sensor types, each with unique characteristics that impact imaging. In recent years, diverse sensors such as pinhole cameras, fisheye cameras [62, 45] with 180° field of views (FoV), and even panoramic cameras [3, 1, 65] with 360° FoV have been increasingly adopted across vision tasks.
@@
Based on our experiments (Fig. 1(a)), we observe that directly applying models designed for narrow-FoV data to wide-FoV data often leads to a suboptimal solution. For instance, SegNext [20] specifically designed for narrow-FoV pinhole cameras typically cannot generalize well when used for analyzing fisheye images with 180° FoV or panoramic images with 360° FoV. This is primarily due to the high degree of image distortion and object deformation present in wide-FoV cameras. As shown in Fig. 2, the altered pixel arrangement in wide-FoV images might lead to misclassifications and limits adaptability of narrow-FoV models [2, 5, 7, 69, 47]. To address this, previous methods have spent a large effort in architecture design to fit specific sensor types, such as CNN-based and transformer-based models [49, 29] for fisheye images, as well as approaches [64, 65] tailored to mitigate distortion in panoramic segmentation.
However, most existing methods deeply couple distortion-aware capability with the entire model (both encoder and decoder). The drawbacks are two-fold: (1) These models are often designed for specific FoVs, yet hard to adapt to varying distortion characteristics. For example, narrow-FoV model SegNext [20] and wide-FoV model 360SFUDA [70] cannot consistently perform well across different wide-FoV segmentation (Fig. 1(a)).
(2) The tight coupling with the overall architecture limits the reusability of large-scale pretrained backbones, as any modification requires retraining. Therefore, in this work, we focus on flexible decoder designs and constrain distortion awareness to the decoder level, yielding greater flexibility in utilizing diverse pretrained backbones, striking a tradeoff between adaptability and modularity.
@@
Figure 2: Comparison of camera imaging with narrow-FoV (left) and wide-FoV (right) cameras. Narrow-FoV cameras maintain geometric fidelity but limited coverage, whereas wide-FoV cameras offer expansive scene capture while introducing substantial geometric distortions.
Furthermore, we mainly utilize the Mamba [17] mechanism due to its benefit of linear computational complexity. However, most previous works [72, 34, 58, 24, 30, 4, 73] have focused on Mamba-based encoders, research on corresponding decoders for downstream tasks remains limited. We contend that decoders are equally crucial as encoders in the context of dense prediction tasks. As illustrated in Fig. 1(b), a widely used decoder UperHead [54] excels at multi-scale feature fusion but suffers from excessive computational overhead. Alternative methods [57, 38] applied lightweight CNN- and transformer-based decoders to reduce computational cost. However, they suffer notable performance degradation when processing wide-FoV data.
To address these challenges, we propose the Deformable Mamba Decoder , an efficient distortion-aware decoder for image segmentation across varying wide-FoVs. In this work, a unified framework is created to effectively handling images from multiple camera types and sources, spanning FoVs from 180° to 360°, from indoor to outdoor scenarios, in synthetic and real-world imageries. Our deformable decoder can seamlessly integrate with CNN-, transformer-, and Mamba-based backbones, imparting distortion-awareness capability to the overall model.
To evaluate the effectiveness of our method, we conduct extensive experiments across five datasets for wide-FoV segmentation, including two 180° (WoodScape [62], SynWoodScape [45]) datasets and three 360° ( Stanford2D3D [1], Matterport3D [3], SynPASS [65]). Our framework performs consistently well across various datasets. Notably, while reducing the decoder parameters by $72\%$ and FLOPs by $97\%$, Deformable Mamba achieved state-of-the-art performance, with a ${+}2.5\%$ absolute improvement over previous VMamba [34] on the 360° Stanford2D3D dataset. To summarize, our contributions are as follows.
  • We propose an efficient decoder based on the Mamba architecture, which addresses the gap in decoder-oriented Mamba-based efficient approaches.
  • For the first time, we incorporate distortion-awareness exclusively at the Mamba decoder level, decoupling it from the overall framework, which enables integration with CNN-, Transformer-, and Mamba-based backbones to enhance distortion perception.
@@ 2.1 Wide Field of View Segmentation
Wide-FoV image segmentation has received increasing attention as it enables broader scene understanding in applications such as autonomous driving and robotics. Wide-FoV images, like panoramic and fisheye images, offer expansive visual coverage but introduce unique challenges due to geometric distortions, spatial resolution variations, and uneven feature distribution. Early segmentation methods [36, 2, 5, 16, 25, 63, 63, 33, 32, 69, 47, 55, 10, 18, 6, 67] have achieved significant success on narrow-FoV image segmentation. However, directly applying pinhole models to wide-FoV images often results in downgraded performance due to object deformations and image distortions. Previous work [59, 60] propose the Panoramic Annular Semantic Segmentation (PASS) framework using a panoramic annular lens system.  [26] introduces panoramic panoptic segmentation for panoramic views. In [64, 65], deformable modules are proposed to enhance the distortion perception capability in transformers. Zheng et al. [68] introduce the Open Panoramic Segmentation (OPS) task in an open-vocabulary context. Apart from panoramas, previous methods [29, 12] revolve around multi-task on fisheye images, while a pretraining paradigm [39] and semi-supervised approaches [42, 40] are introduced to address the challenges of fisheye data. Manzoor et al. [37] substitutes the conventional convolutional modules in U-Net [44] and residual U-Net [43] with restricted deformable convolution (RDC) [12], imparting fisheye-distortion awareness to the model. Departing from these approaches, we are the first to propose a unified model applicable to all wide-FoV inputs, including panoramic and fisheye images.
2.2 Vision Mambas
CNN [28] have achieved notable success in computer vision with efficient local feature extraction and hierarchical learning, while Transformer [51] excel at modeling long-range dependencies despite their quadratic complexity. Recently, the state-space-model (SSM)-based Mamba [17] architecture has emerged as a promising alternative, offering linear complexity and parallel computation through its selective scan mechanism, enabling efficient long-sequence modeling. The application of Mamba in sequence modeling has naturally extended to vision tasks. The serialized processing of input tokens in the Mamba architecture presents notable challenges when transferring to vision tasks. Zhu et al. [72] serialize the image using patch embedding [13], then scan the sequence from two directions. In [34], a quadri-directional scanning mechanism is proposed to supersede Mamba’s initial 1D scan operation to 2D scan. Subsequent work [24, 41, 23, 30, 66, 61, 4] introduce various scanning methods to explore variations of Mamba in visual representations. However, these works have primarily leveraged the efficiency of the Mamba architecture from an encoder perspective, while its application in the decoder remains relatively constrained. Besides, while benefiting from the efficiency of the Mamba architecture, its application in wide-FoV dense prediction tasks remains unexplored. Therefore, for the first time, we propose a mamba-based decoder to address wide-FoV image segmentation.
@@ 3.1 Overall Architecture
Specifically, given the wide-FoV input image of shape $H\times W\times 3$, a stem module is firstly utilized to patchify the original image into the shape of $\frac{H}{4}\times\frac{W}{4}\times 3$, the encoder gradually down-samples feature maps in the $l^{th}$ stage with channel dimensions $C_{l}\in\{96,192,384,768\}$ and resolutions $R_{l}\in\{\frac{H}{4}\times\frac{W}{4},\frac{H}{8}\times\frac{W}{8},\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{H}{32}\}$. These features are hierarchically fused in our proposed Deformable Mamba Decoder. The output features from the decoder’s final layer are first processed through a fully connected layer and subsequently transformed via feature channel convolutions to generate the predicted semantic map.
@@ 3.2 Deformable Mamba Decoder
The Deformable Mamba Decoder comprises four efficient fusion and upsampling modules. Within this framework, we propose a novel Deformable Mamba Fusion (DMF) block, which effectively integrates multi-scale features while enabling the decoder to adaptively process distorted data. Considering the diverse types and degrees of distortion present in different wide-FoV data, our method is not tailored to any specific dataset. Instead, we incorporate distortion awareness into the model in a simple, broadly applicable manner.
As shown in lower part of Fig. 3, the DMF block takes two inputs, $E_{i}$ and $D_{j}$, where $E_{i}$ represents the multi-scale feature maps output by the encoder, and $D_{j}$ denotes the features fused by the DMF block. Here, indices $i$ and $j$ range from 1 to 4 and 0 to 3, respectively, following the constraint $i+j=4$. Specifically, for the first stage of DMF block, both inputs are derived from the feature maps of the last encoder layer, i.e., $E_{4}=D_{0}$. To maximize efficiency by minimizing decoder computational complexity, we employ only a single quadri-directional scanning layer, originally introduced in the 2D selective scan (SS2D) block by Liu et al. [34] for embedding feature $D_{j}$ into the state space.
Figure 4: For an embedded 2D sequence, Mamba [17] employs a uni-directional scan from start to end (1). [72] utilizes a bi-directional scan combining (1) and (2), while [34] adopts a quadri-directional scanning incorporating (1), (2), (3), and (4).
We posit that enabling model distortion-aware capability hinges on providing the model with sufficient deformable capacity, thereby endowing the model with the intrinsic ability to adaptively learn and generalize from distorted data representations. Deformable Convolutions (DCNs) [11, 74, 53, 56] render the rigid kernels in conventional convolution more flexible and dynamically adaptive by introducing data-dependent offsets, which allows the fixed sampling locations to be dynamically adjusted based on the data. [75] provides the Transformer-based model with a data-driven receptive field, enabling the model’s perception of image distortions and deformations. SSM, which describes the dynamics of a system, can be formulated as:
where $\mathbf{x}(\mathbf{t})$ and $\mathbf{y}(\mathbf{t})$ denote the inputs and outputs of the system, respectively, $\mathbf{h}(\mathbf{t})$ is the inner states of the system, $\mathbf{A},\mathbf{B},\mathbf{C}$ are the system parameters.
Mamba [17] and its visual variants [72, 34, 24, 41, 23, 30, 66, 61, 4] are based on SSM. As shown in Fig. 4, while methods such as [72, 4] have attempted to model the spatial information of images by expanding scanning paths, the inherent one-by-one token processing makes it more challenging to integrate a learnable deformation mechanism into state-space modeling, compared to CNN- and Transformer-based architectures. To address this, we adopt a compromise solution by following [65, 64, 68], we incorporate deformable convolution to augment the model’s deformable capacity. In particular, we feed the encoder feature $E_{i}$ into one single DCNv2 [74] layer with a kernel size of $\mathbf{3\times 3}$, which can be mathematically formulated as follows:
$\mathbf{E_{out}}(\mathbf{p})=\sum^{K}_{k=1}\mathbf{w}_{k}\mathbf{m}_{k}\mathbf{E_{in}}(\mathbf{p}+\mathbf{p}_{k}+\Delta\mathbf{p}_{k}),$ (2)
@@
where $\mathbf{p}$ represents the current location in feature map, $\mathbf{w_{k}}$ and $\mathbf{p_{k}}$ denote the weight and initial offsets for the $\mathbf{k}$-th location, respectively. The parameter $\mathbf{m_{k}}$ is a learnable modulation scalar that modulates the input feature amplitudes from different spatial locations. In our implementation, $\mathbf{K}=9$ and $\mathbf{p_{k}}\in{(-1,-1),...,(1,1)}$ defines a $\mathbf{3\times 3}$ convolutional kernel with dilation 1.
The outputs from both branches, each with dimensions $\{h\times w\times c\}$, are concatenated to form a tensor with dimensions $\{h\times w\times 2c\}$, which is subsequently processed by an upsampling module. To fully preserve the feature information from both branches while endowing the model with additional learnable capacity, we refrain from employing parameter-free upsampling methods such as bilinear interporlation. Instead, after an initial feature fusion via a series of $\mathbf{3\times 3}$ convolutions, we enhance the resolution of feature maps using the Pixelshuffle [46] operation, which can be formulated as:
$\mathcal{PS}(T)_{h,w,2c}=T_{2h,2w,\frac{c}{2}},$ (3)
where $\mathcal{PS}$ is an periodic shuffling operator that rearranges the elements of a input tensor $T$ from shape $\{h\times w\times 2c\}$ to $\{2h\times 2w\times\frac{c}{2}\}$.
The upsampled features are linearly projected as the output of the DMF block. Uniquely, the feature map in the final stage, after undergoing convolutional fusion, is not subjected to the PixelShuffle operation for channel reduction and resolution enhancement. It is worth noting that in all the experiments presented in Sec. 4, the decoder architecture is kept consistent, with no dataset-specific modifications. The only adjustment made is to the final output layer, which is tailored to accommodate the varying semantic categories across the datasets.
Our Deformable Mamba Decoder leverages the advantages of the Mamba architecture’s linear computational complexity, resulting in superior computational efficiency. Simultaneously, it addresses the challenge that Mamba, in contrast to CNNs and Transformers, lacks the architectural ability to provide deformation awareness at the feature level. Besides, we introduce distortion awareness exclusively from the decoder perspective, decoupling it from the overall framework and thus enhancing the model’s flexibility and adaptability. Comprehensive qualitative and quantitative analyses in Sec. 4 showcase the exceptional performance and efficiency of our decoder.